# Intelligence Test

### v1

@@ 4.5.2 Mathematics
Figure 7: Results of Intelligence Test on Mathematics. The first two rows test the model on simple addition and subtraction of two-digit numbers, while the last row tests whether the model can provide the solution process for a math competition problem. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results suggest that models are better at reasoning through complex problems than performing simple addition and subtraction.
@@ 4.5.3 Question-Answering
Figure 8: Experimental Results of Intelligence Test in Question-Answering. The three rows correspond to three datasets, and figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models remain at Limited Level.
@@ 4.5.4 Writing
Figure 9: Results of Intelligence Test for Writing in Different Domains. The first row shows language modeling results on Wikipedia. The other two rows present QWen2.5 72B’s writing performance in other domains. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models are at Limited Level.
@@
Figure 10: Results of Intelligence Test for Writing in Different Languages. The first row shows language modeling results in Polish. The other two rows present QWen2.5 72B’s writing performance in different languages. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results show that all models lie withat Limited Level.
@@ 4.6 Revisiting Current AI Techniques
In previous subsections, we evaluate current AI systems in areas such as vision, search, recommendation, and language. We can see that AI remains at the Limited Level. Although this insight was not widely recognized before this study, we will show that it has profoundly impacted existing AI technologies. In other words, current AI technologies are exactly developed in the context of Limited-Level intelligence.
@@
We begin by establishing a connection between AI technology and Intelligence Test through the concept of loss. Loss plays a crucial role in AI, especially in deep learning, as it quantifies the degree of error made by an AI system. The smaller the loss, the more advanced the AI system is considered. We can see that the concept of loss is very similar to the concept of failure count in Intelligence Test, where failure count quantifies the extent of subjects’ errors. The smaller the failure count, the more intelligent the subject is considered. Therefore, we can treat failure count as a form of loss, which we will refer to as “Intelligence Test Loss” in this subsection.
@@
Intelligence Test Loss has a strong physical meaning and naturally reflects the performance of AI systems. Yet, most advanced AI systems are stuck at the Limited Level and Intelligence Test Loss diverges, making directly adopting this loss infeasible. In the following, we demonstrate that many current AI technologies are profoundly related to the divergence of Intelligence Test Loss, even though these technologies were not explicitly designed or used with this awareness in mind.
@@
Hard Negative Sampling: Hard negative sampling is a widely used optimization technique in many AI fields, including vision (Shrivastava et al., 2016), search (Zhan et al., 2021), recommendation (Ding et al., 2020), and language (Kalantidis et al., 2020) tasks. It penalizes the model’s top-k most incorrect predictions (i.e., hard negatives), rather than punishing all of its wrong predictions (i.e., random negatives). Researchers explain its effectiveness with various hypotheses, such as increasing gradient magnitudes (Xiong et al., 2021), bootstrapping the training data (Shrivastava et al., 2016), simulating an easy-to-hard curriculum learning process (Chen et al., 2021b), etc. However, from the perspective of Intelligence Test Loss, its effectiveness becomes easy to understand. Since Hard Negative Sampling focuses on top-$k$ errors, its loss can be seen as $\min(\text{Intelligence Test Loss},k)$. By truncating Intelligence Test Loss with $k$, this approach ensures convergence. This truncation operation gives up on the difficult cases where models fail more than $k$ times before finding the correct solutions. It only optimizes performance in easy cases where the failure count can be smaller than $k$. Ignoring poor performance in difficult cases enables the model to focus on improving accuracy in simple cases. This approach is effective for Limited-Level intelligence, but if the model could reach the Autonomous Level, there would be no need to ignore difficult cases, and this method would not be so effective.
@@
Cross-Entropy Loss: Cross-entropy loss is a commonly used loss function, widely applied across tasks such as vision (Oord et al., 2018; Radford et al., 2021) and language (Radford et al., 2019; Izacard et al., 2021). Earlier researchers provided heuristic explanations for its effectiveness, such as making the neural network’s embedding distribution more uniform (Wang & Isola, 2020) and automatically weighting different negatives (Chen et al., 2020). Yet, from the perspective of Intelligence Test Loss, its role becomes clearer. It can be seen as using $\log(\text{Intelligence Test Loss})$ as the loss. $\log$ transformation leads to better convergence. For instance, if Intelligence Test Loss follows a power-law distribution with an exponent between 1 and 2, its expectation does not converge, but the $\log$ transformation does. In this way, Cross-Entropy loss helps address the divergence of Intelligence Test Loss and thus makes the training process more effective. Nevertheless, if models were at the Autonomous Level, it would not be so effective since the loss would already be convergent.
Reinforcement Learning (RL): RL (Kaelbling et al., 1996) lets AI systems explore solutions themselves and rewards them when they succeed. It is similar to how animals learn. However, its application is limited because the training cost is prohibitively high (Dulac-Arnold et al., 2021). We can explain this high training cost based on the convergence of Intelligence Test Loss at Limited Level. The cost of RL is closely related to the number of failed attempts, which is indeed Intelligence Test Loss. Therefore, the cost is infinite at Limited Level, making RL infeasible. Recently, DeepSeek-R1 (Guo et al., 2025) shows that RL can be applied in mathematical and coding tasks. This is because current advanced models are approaching the Capable Level in the two areas, as reflected by our previous experiments in Section 4.5.1 and 4.5.2. Models at Capable Level are more likely to find correct answers and their cost in the RL process is much lower. Yet for many other tasks, such as writing, current models are still far from the Capable Level, making RL difficult to apply.
We can see that Intelligence Test provides deep insights for understanding AI techniques. Although earlier researchers mainly designed algorithms based on heuristics without the knowledge of this test, Intelligence Test effectively reveals the fundamental reasons behind their success. With the guidance of Intelligence Test, we believe researchers will easily design more advanced AI techniques in the future.
5 Future Prediction by Intelligence Test
In this section, we make predictions about the model size needed to achieve the Autonomous Level in language tasks. We begin by introducing the empirical relationship between model size and decay rate in Intelligence Test. Using this relationship, we will extrapolate to even larger model sizes, making predictions about the future trajectory of AI development.
@@ 5.1 Fitting Failure Count Distribution
In this subsection, we aim to quantitatively characterize the performance of different models in Intelligence Test. Take a look at the experimental results from the previous section, such as Figure 9 and 10. The distribution of failure count is close to a straight line in a log-log plot, especially for failure count between $10$ and $100$. When the failure count exceeds $100$, the data points become scattered. When the failure count is less than $10$, the distribution is a curve that bends downward with an increasingly steeper slope. Since a straight line in a log-log plot suggests a power law distribution, this observation suggests that the failure count distribution can be approximated by a power law, especially when the failure count is neither too small nor too large.
@@ 5.3 Predicting Future
6 Theoretical Analysis of Intelligence Test
@@
To address this, we turn to the framework of Self-Organized Criticality (SOC) (Bak et al., 1987), a complexity theory in Physics. SOC describes systems in a criticality state where small perturbations can trigger large-scale changes. We propose that Intelligence Test is deeply related to SOC and that drawing this connection provides insights into the nature of human tasks and AI. In the following, we will first discuss how Intelligence Test relates to SOC, then model Intelligence Test with a SOC framework.
@@ 6.1 Human Tasks exhibit Criticality
We propose that human tasks exhibit criticality, which leads to Intelligence Test as a SOC system. In Intelligence Test, the questions posed and the corresponding correct answers form a system. The self-organizing nature of this system is evident in the way that different questions correspond to different correct answers. We can imagine the correspondence between a question and its answer as a dynamic process. When a question and the correct answer are given, the system is in a stable state. When certain parts of the question are modified, the answer should undergo a self-organizing modification process and eventually evolve into a new correct answer, thus bringing the entire system back to stability. The system’s criticality arises from the nature of human tasks. For example, a minor alteration in a mathematical question can drastically change the approach required to solve it. This criticality property requires the test subject to address subtle differences in the question that can lead to significantly distinct answers. Merely memorizing answers for several specific cases does not help because a small change results in entirely different answers. We believe this is the cause of why so many AI systems are at Limited Level. AI systems might simply memorize some answers and yet human tasks are not friendly to memorization because of criticality property.
@@
Physics: Physics exhibits criticality. In physics, even seemingly similar problems can lead to vastly different solutions depending on the initial conditions. For instance, consider a question about a matter’s state or superconductivity. The answer relies crucially on whether the temperature is above or below a threshold. Similarly, if we ask about physical laws, the appropriate theory, quantum mechanics or classical physics, depends on the scale. This phenomenon is ubiquitous in physics: small variations in initial conditions can lead to fundamentally different results. By incorporating such questions into an Intelligence Test, we naturally create a system with SOC property. The test subjects must adapt to subtle changes in initial conditions. Otherwise, their responses would be entirely wrong. If a participant relies solely on memorizing answers for some specific conditions, they will struggle to apply their knowledge to new, subtly altered questions. In such cases, the failure attempts are likely to approach infinity, as the answers will deviate drastically from the memorized solutions.
Mathematics: Mathematics exhibits criticality. Mathematics embodies the very essence of SOC. Take Fermat’s Last Theorem as an example. It asserts that there are no integer solutions to the equation $x^{n}+y^{n}=z^{n}$ for any integer $n>2$. At first glance, the change of $n$ might seem like a minor adjustment and does not affect the essence of the problem. However, the theorem’s sensitivity to the value of n is profound. As soon as $n$ grows larger, the problem becomes much more complex. The theorem with small $n$ values was proven soon, but it took more than two centuries before the general case was finally proven. Many math problems show such criticality property that seemingly small modifications can alter the entire landscape of solutions. Thus, when mathematical problems are used in Intelligence Test, subjects must be acutely aware of the details and select the appropriate mathematical tools. Subjects that do not possess this ability will take infinite attempts to arrive at the correct answer.
Law: Legal issues exhibit criticality. In law, seemingly minor differences in behavior can lead to vastly divergent legal consequences. For instance, a suspect’s actions might determine whether the charge is premeditated murder or voluntary manslaughter, whether it was excessive self-defense or justifiable defense, or even whether they are guilty or innocent. Such significant shifts in legal outcomes can arise from differences in behavior that, on the surface, might appear negligible. When legal questions are incorporated into an Intelligence Test, they require the subjects to discern these nuances and reason the outcomes. If a subject has only memorized conclusions for specific cases, they will struggle in new scenarios and will inevitably make infinite attempts before arriving at the right answers.
@@
These human tasks exhibit criticality and make Intelligence Test exhibit SOC. Criticality is a typical symbol of complexity: Since a slight change in the question can result in an entirely new answer, successfully operating the task requires a full understanding of the underlying mechanism. Otherwise, subjects will be completely wrong about the correct answer. It will be difficult for them to arrive at the right one through trial and error. This is indeed the case for current AI systems as shown in our experiments. They remain at the Limited Level where the number of trials is infinite. This suggests that most AI systems do not fully understand the mechanism. Even if they are trained on enormous data and have memorized all of it, it is not enough for tasks with criticality properties.
6.2 Modeling Intelligence Test as a Self-Organized Criticality System
@@
We have qualitatively explored how Intelligence Test exhibits SOC property on human tasks. Now, we will adopt a quantitative perspective to validate this hypothesis. We will see that it closely resembles a typical SOC model and exhibits very similar experimental phenomena.
@@
Figure 14: Modeling Intelligence Test as a complex network. Nodes represent concepts and edges represent interdependence. The network is also a sandpile model (Bak et al., 1987).
@@
In Intelligence Test, both questions and answers inherently consist of many basic concepts. Since answers change when questions change, these concepts are intricately coupled. Because of the criticality property of human tasks, even a small change in one concept can trigger significant changes in many other concepts. We can imagine this interconnection as a network graph, as shown in Figure 14. Each node represents a concept, and the edges between the nodes represent the couplings between these concepts. Slightly changing the question is akin to disturbing a node. When a node becomes unstable, it is activated and is likely to make its neighbors also unstable. This might eventually result in cascading activation throughout the network. Such changes throughout the entire network imitate how seemingly small changes in the question can lead to very different answers in Intelligence Test.
@@
This conceptual network is very similar to a typical SOC system, namely the sandpile model (Bak et al., 1987). It has been used to model complex systems in the natural world. In the sandpile model, each node accumulates sand. To disturb the network is to add a little more sand on a node. When the amount of sand exceeds a certain threshold, the sand topples from this node. The toppled sand is distributed to its neighbors, and thus the neighbors’ sand may also topple. This cascade of toppling sometimes results in an avalanche throughout the entire network. We believe that this sand avalanche effect is close to the cascading activation of concepts in our conceptual network for Intelligence Test. Therefore, we further implement our conceptual network as a sandpile model. Specifically, a slight change in a question is equivalent to adding a small amount of sand to a node. The degree to which the answer changes after altering the question is analogous to the avalanche size triggered by adding sand. Through this correspondence, we model Intelligence Test as a sandpile model. The validity of this modeling can be confirmed by testing whether the sandpile model accurately reflects the characteristics of Intelligence Test as observed in real-world applications.
@@
Modeling Intelligence Test as a sandpile model can explain why the failure count distribution of current AI systems resembles a power law. In real-world experiments, we empirically show that failure count distribution is very close to a power law, as described in Section 5.1. However, the underlying reason is a mystery. Now, we can investigate this phenomenon based on our abstraction of Intelligence Test, namely the sandpile model. Figure 15 shows the simulated distribution of avalanche size in the sandpile model. Based on our analogy, this also corresponds to the distribution of how much answers change after altering the questions in Intelligence Test. The distribution follows a power law, which closely matches the failure count distribution of AI systems. This actually reflects that current AI systems rely on memorization and exploration to solve new problems. Here is the reason. If AI relies on memorization and exploration, it means that when faced with a new question in Intelligence Test, it recalls similar questions it has memorized and, based on the remembered answers, explores potential solutions. As a result, the number of explorations correlates with the distance between the answers. In contrast, if the AI system genuinely understood the test questions, the distance between answers would not correlate with the number of failure attempts. In fact, the mechanics of the sandpile model are clear, and an AI system that truly grasps these mechanics could directly compute the stable state without any failure attempt. Therefore, the phenomena indicate that AI systems do not fully understand the mechanics of the task and instead rely on memorization and exploration to find answers.
@@
Modeling Intelligence Test as a sandpile model also helps to illuminate how scaling works. From our experiments in Section 5.2, we empirically observe that scaling model size can improve the decay rate and thereby improve intelligence. Yet, the underlying reason is a mystery. However, we find that the dimensionality of the sandpile topology has a similar effect and can explain this phenomenon. More precisely, both theoretical analyses (Dhar, 2006; Zachariou et al., 2015) and our simulations in Figure 15 show that the power law exponent of the sandpile model increases as the topology becomes high-dimensional. This is because a higher-dimensional topology creates more connections between nodes, enabling the network to stabilize more quickly when disturbed. We can draw the analogy between the sandpile’s stabilization through these connections and intelligence systems’ solving new problems through exploration. When we scale an intelligent system, we expand the problem-solving space to a higher-dimensional level and construct new paths between concept nodes. These new paths enable current intelligence systems to explore new solutions from memorized ones more quickly. This is akin to how a sandpile model stabilizes itself through new connections in a higher dimensionality. If we regard these new paths for exploration as the new connections in SOC systems, the effects of scaling can be explained: scaling makes exploration more effective.
In summary, Intelligence Test can be modeled as a sandpile model. It helps us gain a deep understanding of the nature of human tasks and current AI. The sandpile model is a complex system where a little more sand can result in an avalanche throughout the entire landscape. Similarly, Intelligence Test with human tasks are also complex where a small change in environments requires completely different responses. No matter whether it is to predict the state of the sandpile model or find the correct solution in Intelligence Test, it is necessary to fully understand the underlying mechanisms. Yet, most AI systems do not understand these mechanisms and rely on superficial imitation, such as memorization and exploration. They exhibit a typical power law with a small exponent. Scaling is able to transform the exploration space to a higher dimensionality and makes the exploration process more effective. Nevertheless, if AI lacks a full understanding of underlying mechanisms, SOC property will make it extremely difficult to achieve the Autonomous Level of intelligence.

### v2

@@ 4.5.2 Mathematics
Figure 8: Results of Survival Game on Mathematics. The first two rows test the model on simple addition and subtraction of two-digit numbers, while the last row tests whether the model can provide the solution process for a math competition problem. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results suggest that models are better at reasoning through complex problems than performing simple addition and subtraction.
@@ 4.5.3 Question-Answering
Figure 9: Experimental Results of Survival Game in Question-Answering. The three rows correspond to three datasets, and figures in different columns correspond to different models. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models remain at Limited Level.
@@ 4.5.4 Writing
Figure 10: Results of Survival Game for Writing in Different Domains. The first row shows language modeling results on Wikipedia. The other two rows present QWen2.5 72B’s writing performance in other domains. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results indicate that all models are at Limited Level.
@@
Figure 11: Results of Survival Game for Writing in Different Languages. The first row shows language modeling results in Polish. The other two rows present QWen2.5 72B’s writing performance in different languages. The red line is a power law curve drawn based on the distribution of the model’s data points. Its exponent roughly represents the model’s failure decay rate. Results show that all models lie withat Limited Level.
@@ 4.6 Revisiting Current AI Techniques
In previous subsections, we evaluate current AI systems in areas such as vision, search, recommendation, and language. We can see that AI remains at the Limited Level. Although this insight was not widely recognized before this study, we find that it has already profoundly impacted existing AI technologies. In other words, current AI technologies are exactly developed in the context of Limited-Level intelligence.
@@
We begin by establishing a connection between AI technology and Survival Game through the concept of loss. Loss plays a crucial role in AI, especially in deep learning, as it quantifies the degree of error made by an AI system. The smaller the loss, the more advanced the AI system is considered. We can see that the concept of loss is very similar to the concept of failure count in Survival Game, where failure count quantifies the extent of subjects’ errors. The smaller the failure count, the more intelligent the subject is considered. Therefore, we can treat failure count as a form of loss, which we will refer to as “Survival Game Loss” in this subsection.
@@
Survival Game Loss has a strong physical meaning and naturally reflects the performance of AI systems. Yet, most advanced AI systems are stuck at the Limited Level and Survival Game Loss diverges, making directly adopting this loss infeasible. In the following, we demonstrate that many current AI technologies are profoundly related to the divergence of Survival Game Loss, even though these technologies were not explicitly designed or used with this awareness in mind.
@@
Hard Negative Sampling: Hard negative sampling is a widely used optimization technique in many AI fields, including vision (Shrivastava et al., 2016), search (Zhan et al., 2021), recommendation (Ding et al., 2020), and language (Kalantidis et al., 2020) tasks. It penalizes the model’s top-k most incorrect predictions (i.e., hard negatives), rather than punishing all of its wrong predictions (i.e., random negatives). Researchers explain its effectiveness with various hypotheses, such as increasing gradient magnitudes (Xiong et al., 2021), bootstrapping the training data (Shrivastava et al., 2016), simulating an easy-to-hard curriculum learning process (Chen et al., 2021b), etc. However, from the perspective of Survival Game Loss, its effectiveness becomes easy to understand. Since Hard Negative Sampling focuses on top-$k$ errors, its loss can be seen as $\min(\text{Survival Game Loss},k)$. By truncating Survival Game Loss with $k$, this approach ensures convergence. This truncation operation gives up on the difficult cases where models fail more than $k$ times before finding the correct solutions. It only optimizes performance in easy cases where the failure count can be smaller than $k$. Ignoring poor performance in difficult cases enables the model to focus on improving accuracy in simple cases. This approach is effective for Limited-Level intelligence, but if the model could reach the Autonomous Level, there would be no need to ignore difficult cases, and this method would not be so effective.
@@
Cross-Entropy Loss: Cross-entropy loss is a commonly used loss function, widely applied across tasks such as vision (Oord et al., 2018; Radford et al., 2021) and language (Radford et al., 2019; Izacard et al., 2021). Earlier researchers provided heuristic explanations for its effectiveness, such as making the neural network’s embedding distribution more uniform (Wang & Isola, 2020) and automatically weighting different negatives (Chen et al., 2020). Yet, from the perspective of Survival Game Loss, its role becomes clearer. It can be seen as using $\log(\text{Survival Game Loss})$ as the loss. $\log$ transformation leads to better convergence. For instance, if Survival Game Loss follows a power-law distribution with an exponent between 1 and 2, its expectation does not converge, but the $\log$ transformation does. In this way, Cross-Entropy loss helps address the divergence of Survival Game Loss and thus makes the training process more effective. Nevertheless, if models were at the Autonomous Level, it would not be so effective since the loss would already be convergent.
Reinforcement Learning (RL): RL (Kaelbling et al., 1996) lets AI systems explore solutions themselves and rewards them when they succeed. It is similar to how animals learn. However, its application is limited because the training cost is prohibitively high (Dulac-Arnold et al., 2021). We can explain this high training cost based on the convergence of Survival Game Loss at Limited Level. The cost of RL is closely related to the number of failed attempts, which is indeed Survival Game Loss. Therefore, the cost is infinite at Limited Level, making RL infeasible. Recently, DeepSeek-R1 (Guo et al., 2025) shows that RL can be applied in mathematical and coding tasks. This is because current advanced models are approaching the Capable Level in the two areas, as reflected by our previous experiments in Section 4.5.1 and 4.5.2. Models at Capable Level are more likely to find correct answers and their cost in the RL process is much lower. Yet for many other tasks, such as writing, current models are still far from the Capable Level, making RL difficult to apply.
We can see that Survival Game provides deep insights for understanding AI techniques. Although earlier researchers mainly designed algorithms based on heuristics without the knowledge of this test, Survival Game effectively reveals the fundamental reasons behind their success. With the guidance of Survival Game, we believe researchers will easily design more advanced AI techniques in the future.
5 Scaling in Survival Game
In this section, we make predictions about the model size needed to achieve the Autonomous Level in language tasks. We begin by introducing the empirical relationship between model size and decay rate in Survival Game. Using this relationship, we will extrapolate to even larger model sizes, making predictions about the future trajectory of AI development.
@@ 5.1 Fitting Failure Count Distribution
In this subsection, we aim to quantitatively characterize the performance of different models in Survival Game. Take a look at the experimental results from the previous section, such as Figure 10 and 11. The distribution of failure count is close to a straight line in a log-log plot, especially for failure count between $10$ and $100$. When the failure count exceeds $100$, the data points become scattered. When the failure count is less than $10$, the distribution is a curve that bends downward with an increasingly steeper slope. Since a straight line in a log-log plot suggests a power law distribution, this observation suggests that the failure count distribution can be approximated by a power law, especially when the failure count is neither too small nor too large.
@@ 5.2 Fitting Effect of Scaling
Besides the results of QWen 2.5, we also show the performance across different model architectures in Figure 1. The models include GPT2 (Radford et al., 2019), OPT (Zhang et al., 2022), GPT-Neo (Black et al., 2021), Llama-1 (Touvron et al., 2023a), Llama-2 (Touvron et al., 2023b), Llama-3 (Dubey et al., 2024), Phi-2 (Javaheripi et al., 2023), GLM4 (GLM et al., 2024), DeepSeek-V2 (DeepSeek-AI, 2024), BaichuanM1 (Baichuan, 2024), Mistral (Jiang et al., 2023), QWen2.5 (Yang et al., 2024b). We can see that across different model architectures, larger models tend to perform better, forming an approximately linear trend. This is evident in models such as the OPT series, Llama 3 & 3.2 series. This observation aligns with our idea of fitting a straight line. Additionally, we observe that more recently released models tend to achieve better performance over time. Therefore, optimistically speaking, although the fitted line slightly overestimates the performance of large models, their performance are expected to improve over time. Moreover, we can see that among these models, Qwen2.5 demonstrates state-of-the-art performance. As a result, we use Qwen2.5 as a benchmark to represent the current cutting-edge level.
@@ 5.3 Predicting Future
6 Theoretical Analysis of Survival Game
@@
To address this, we turn to the framework of Self-Organized Criticality (SOC) (Bak et al., 1987), a complexity theory in Physics. SOC describes systems in a criticality state where small perturbations can trigger large-scale changes. We propose that Survival Game is deeply related to SOC and that drawing this connection provides insights into the nature of human tasks and AI. In the following, we will first discuss how Survival Game relates to SOC, then model Survival Game with a SOC framework.
@@ 6.1 Human Tasks exhibit Criticality
We propose that human tasks exhibit criticality, which leads to Survival Game as a SOC system. In Survival Game, the questions posed and the corresponding correct answers form a system. The self-organizing nature of this system is evident in the way that different questions correspond to different correct answers. We can imagine the correspondence between a question and its answer as a dynamic process. When a question and the correct answer are given, the system is in a stable state. When certain parts of the question are modified, the answer should undergo a self-organizing modification process and eventually evolve into a new correct answer, thus bringing the entire system back to stability. The system’s criticality arises from the nature of human tasks. For example, a minor alteration in a mathematical question can drastically change the approach required to solve it. This criticality property requires the test subject to address subtle differences in the question that can lead to significantly distinct answers. Merely memorizing answers for several specific cases does not help because a small change results in entirely different answers. We believe this is the cause of why so many AI systems are at Limited Level. AI systems might simply memorize some answers and yet human tasks are not friendly to memorization because of criticality property.
@@
Physics: Physics exhibits criticality. In physics, even seemingly similar problems can lead to vastly different solutions depending on the initial conditions. For instance, consider a question about a matter’s state or superconductivity. The answer relies crucially on whether the temperature is above or below a threshold. Similarly, if we ask about physical laws, the appropriate theory, quantum mechanics or classical physics, depends on the scale. This phenomenon is ubiquitous in physics: small variations in initial conditions can lead to fundamentally different results. By incorporating such questions into an Survival Game, we naturally create a system with SOC property. The test subjects must adapt to subtle changes in initial conditions. Otherwise, their responses would be entirely wrong. If a participant relies solely on memorizing answers for some specific conditions, they will struggle to apply their knowledge to new, subtly altered questions. In such cases, the failure attempts are likely to approach infinity, as the answers will deviate drastically from the memorized solutions.
Mathematics: Mathematics exhibits criticality. Mathematics embodies the very essence of SOC. Take Fermat’s Last Theorem as an example. It asserts that there are no integer solutions to the equation $x^{n}+y^{n}=z^{n}$ for any integer $n>2$. At first glance, the change of $n$ might seem like a minor adjustment and does not affect the essence of the problem. However, the theorem’s sensitivity to the value of n is profound. As soon as $n$ grows larger, the problem becomes much more complex. The theorem with small $n$ values was proven soon, but it took more than two centuries before the general case was finally proven. Many math problems show such criticality property that seemingly small modifications can alter the entire landscape of solutions. Thus, when mathematical problems are used in Survival Game, subjects must be acutely aware of the details and select the appropriate mathematical tools. Subjects that do not possess this ability will take infinite attempts to arrive at the correct answer.
Law: Legal issues exhibit criticality. In law, seemingly minor differences in behavior can lead to vastly divergent legal consequences. For instance, a suspect’s actions might determine whether the charge is premeditated murder or voluntary manslaughter, whether it was excessive self-defense or justifiable defense, or even whether they are guilty or innocent. Such significant shifts in legal outcomes can arise from differences in behavior that, on the surface, might appear negligible. When legal questions are incorporated into an Survival Game, they require the subjects to discern these nuances and reason the outcomes. If a subject has only memorized conclusions for specific cases, they will struggle in new scenarios and will inevitably make infinite attempts before arriving at the right answers.
@@
These human tasks exhibit criticality and make Survival Game exhibit SOC. Criticality is a typical symbol of complexity: Since a slight change in the question can result in an entirely new answer, successfully operating the task requires a full understanding of the underlying mechanism. Otherwise, subjects will be completely wrong about the correct answer. It will be difficult for them to arrive at the right one through trial and error. This is indeed the case for current AI systems as shown in our experiments. They remain at the Limited Level where the number of trials is infinite. This suggests that most AI systems do not fully understand the mechanism. Even if they are trained on enormous data and have memorized all of it, it is not enough for tasks with criticality properties.
6.2 Modeling Survival Game as a Self-Organized Criticality System
@@
We have qualitatively explored how Survival Game exhibits SOC property on human tasks. Now, we will adopt a quantitative perspective to validate this hypothesis. We will see that it closely resembles a typical SOC model and exhibits very similar experimental phenomena.
@@
Figure 15: Modeling Survival Game as a complex network. Nodes represent concepts and edges represent interdependence. The network is also a sandpile model (Bak et al., 1987).
@@
In Survival Game, both questions and answers inherently consist of many basic concepts. Since answers change when questions change, these concepts are intricately coupled. Because of the criticality property of human tasks, even a small change in one concept can trigger significant changes in many other concepts. We can imagine this interconnection as a network graph, as shown in Figure 15. Each node represents a concept, and the edges between the nodes represent the couplings between these concepts. Slightly changing the question is akin to disturbing a node. When a node becomes unstable, it is activated and is likely to make its neighbors also unstable. This might eventually result in cascading activation throughout the network. Such changes throughout the entire network imitate how seemingly small changes in the question can lead to very different answers in Survival Game.
@@
This conceptual network is very similar to a typical SOC system, namely the sandpile model (Bak et al., 1987). It has been used to model complex systems in the natural world. In the sandpile model, each node accumulates sand. To disturb the network is to add a little more sand on a node. When the amount of sand exceeds a certain threshold, the sand topples from this node. The toppled sand is distributed to its neighbors, and thus the neighbors’ sand may also topple. This cascade of toppling sometimes results in an avalanche throughout the entire network. We believe that this sand avalanche effect is close to the cascading activation of concepts in our conceptual network for Survival Game. Therefore, we further implement our conceptual network as a sandpile model. Specifically, a slight change in a question is equivalent to adding a small amount of sand to a node. The degree to which the answer changes after altering the question is analogous to the avalanche size triggered by adding sand. Through this correspondence, we model Survival Game as a sandpile model. The validity of this modeling can be confirmed by testing whether the sandpile model accurately reflects the characteristics of Survival Game as observed in real-world applications.
@@
Modeling Survival Game as a sandpile model can explain why the failure count distribution of current AI systems resembles a power law. In real-world experiments, we empirically show that failure count distribution is very close to a power law, as described in Section 5.1. However, the underlying reason is a mystery. Now, we can investigate this phenomenon based on our abstraction of Survival Game, namely the sandpile model. Figure 16 shows the simulated distribution of avalanche size in the sandpile model. Based on our analogy, this also corresponds to the distribution of how much answers change after altering the questions in Survival Game. The distribution follows a power law, which closely matches the failure count distribution of AI systems. This actually reflects that current AI systems rely on memorization and exploration to solve new problems. Here is the reason. If AI relies on memorization and exploration, it means that when faced with a new question in Survival Game, it recalls similar questions it has memorized and, based on the remembered answers, explores potential solutions. As a result, the number of explorations correlates with the distance between the answers. In contrast, if the AI system genuinely understood the test questions, the distance between answers would not correlate with the number of failure attempts. In fact, the mechanics of the sandpile model are clear, and an AI system that truly grasps these mechanics could directly compute the stable state without any failure attempt. Therefore, the phenomena indicate that AI systems do not fully understand the mechanics of the task and instead rely on memorization and exploration to find answers.
@@
Modeling Survival Game as a sandpile model also helps to illuminate how scaling works. From our experiments in Section 5.2, we empirically observe that scaling model size can improve the decay rate and thereby improve intelligence. Yet, the underlying reason is a mystery. However, we find that the dimensionality of the sandpile topology has a similar effect and can explain this phenomenon. More precisely, both theoretical analyses (Dhar, 2006; Zachariou et al., 2015) and our simulations in Figure 16 show that the power law exponent of the sandpile model increases as the topology becomes high-dimensional. This is because a higher-dimensional topology creates more connections between nodes, enabling the network to stabilize more quickly when disturbed. We can draw the analogy between the sandpile’s stabilization through these connections and intelligence systems’ solving new problems through exploration. When we scale an intelligent system, we expand the problem-solving space to a higher-dimensional level and construct new paths between concept nodes. These new paths enable current intelligence systems to explore new solutions from memorized ones more quickly. This is akin to how a sandpile model stabilizes itself through new connections in a higher dimensionality. If we regard these new paths for exploration as the new connections in SOC systems, the effects of scaling can be explained: scaling makes exploration more effective.
In summary, Survival Game can be modeled as a sandpile model. It helps us gain a deep understanding of the nature of human tasks and current AI. The sandpile model is a complex system where a little more sand can result in an avalanche throughout the entire landscape. Similarly, Survival Game with human tasks are also complex where a small change in environments requires completely different responses. No matter whether it is to predict the state of the sandpile model or find the correct solution in Survival Game, it is necessary to fully understand the underlying mechanisms. Yet, most AI systems do not understand these mechanisms and rely on superficial imitation, such as memorization and exploration. They exhibit a typical power law with a small exponent. Scaling is able to transform the exploration space to a higher dimensionality and makes the exploration process more effective. Nevertheless, if AI lacks a full understanding of underlying mechanisms, SOC property will make it extremely difficult to achieve the Autonomous Level of intelligence.